# TableMaster: A Recipe to Advance Table Understanding with Language Models

### v3

@@ 5.3 Ablation Study
Reasoning. The reasoning stage exhibits the most significant performance drop when removed. Removing textual reasoning leads to the largest accuracy decline, with a 4.28% drop on WikiTQ and 1.73% on TabFact, underscoring its necessity for complex reasoning tasks. Similarly, removing symbolic reasoning results in a 2.03% and 0.79% drop on WikiTQ and TabFact, respectively, demonstrating that symbolic reasoning enhances numerical and structured table interpretations. Finally, removing textual guidance, which improves the semantic flexibility of symbolic reasoning, reduces accuracy by 2.92% on WikiTQ and 0.44% on TabFact. This highlights that textual guidance is particularly beneficial and important in symbolic reasoning by ensuring alignment with the problem context.
These results reveal that TableMaster ’s performance relies on a combination of structural understanding, content adaptation, and an effective reasoning strategy. Among these, structural extraction and reasoning components contribute the most to performance, with textual reasoning being the most critical. Additionally, while content understanding components such as re-construction and verbalization provide meaningful improvements, their impact varies across datasets. These findings validate the effectiveness of our framework in handling diverse table understanding tasks. More analysis of adaptive reasoning can be found at Appendix L.
@@ 6 Conclusion
In this paper, we explore table understanding with language models. Given the characteristics of tabular data, we identify key challenges in table understanding. To overcome these challenges, we propose TableMaster, a recipe and comprehensive framework that integrates multiple solutions. Extensive analyses and experiments demonstrate our findings and the effectiveness of TableMaster. In the future, we plan to extend and refine the framework to improve its performance across diverse practical applications, where discussed in Appendix C.
Appendix A Ethics and Broader Impacts
This work focuses on improving the table understanding capabilities of large language models through modular reasoning strategies. It does not involve the use of private, sensitive, or human-subject data. All datasets used are publicly available and widely adopted in prior work. The proposed method, TableMaster, can benefit a wide range of practical applications such as web search, spreadsheet automation, and enterprise data querying. We acknowledge that improved table reasoning systems could be misused in generating misleading summaries or automating decision-making without human oversight. However, we expect the modular and transparent nature of our method to encourage responsible deployment. No foreseeable negative societal impacts arise directly from this research.
Appendix B Scope, Key Contributions, and Comparison with Prior Work
Below we outline how TableMaster advances table understanding and how it fundamentally departs from earlier work.
A Recipe and a unifying framework, not a single algorithm.
TableMaster is proposed as a Recipe for general table understanding with language models. While most previous studies focus on a single aspect—often with specialized, heavyweight components—TableMaster supplies four lightweight yet complementary modules that together span the entire processing pipeline. This design makes TableMaster easy to adapt across diverse downstream tasks.
Challenge–solution mapping.
Section 3 and much of the appendix distill four intrinsic properties of tabular data—Structured, Intensive, Concise, Numerical—into four corresponding challenges:
  1. 1.
    Target-data localization
  2. 2.
    Semantic deficiency
  3. 3.
    Numerical inaccuracy in textual reasoning
  4. 4.
    Semantic rigidity in symbolic reasoning
For each challenge we propose a dedicated, minimal solution, whereas earlier efforts typically address only one point and overlook the others.
Efficient subtable extraction and symbolic reasoning.
Prior systems rely on elaborate heuristics for subtable extraction, which often lose information. TableMaster instead couples LLM-based column selection with SQL-style row filtering, then reconstructs a compact Table-of-Focus that preserves context—a simpler yet more general strategy.
Table verbalization for semantic enrichment.
To mitigate semantic deficiency, TableMaster inserts a Table-to-Text verbalization step that enriches input when header cues are sparse. This re-purposes a classic task to bolster table comprehension and, to our knowledge, has not been explored in prior work.
Adaptive textual–symbolic reasoning.
Although chain-of-thought reasoning has improved substantially, most earlier approaches still rely mainly on symbolic logic. We analyze the strengths and weaknesses of both paradigms (Appendix L) and introduce an adaptive scheme that dynamically chooses between—or blends—textual and symbolic reasoning without costly sampling or majority voting.
Impact.
TableMaster achieves consistent gains across five benchmarks, offers an extensible baseline for real-world deployment, and paves the way for next-generation language models equipped for robust, general-purpose table understanding.
Appendix C Limitations, Extendability, and Future Works
@@
C.1 Technical Refinement
Wild Table. In our experiments, the tables in the three datasets we use are already cleaned; therefore, we do not explicitly implement table normalization in our evaluation experiments. However, we conduct analysis experiments to highlight the importance of table normalization for handling wild tables. In practical scenarios, various tools are available for table normalization. Regular expression matching can be employed for formatting, and small language models can also be leveraged to efficiently process and normalize tables [63].
@@ C.1 Technical Refinement
Table-of-Focus Construction. In designing the Table-of-Focus, we employ two efficient methods: LM prompting for column lookup and SQL generation for row lookup. The Table-of-Focus is then constructed based on the results of these two lookups. Many previous works [45, 17] have introduced complex approaches for extracting relevant sub-tables. In contrast, our method remains intentionally simple, prioritizing efficiency and adaptability. We believe that in the future, more advanced techniques may emerge to further enhance the extraction of key information.
Table Verbalization. To facilitate the implementation of TableMaster, we utilize language models themselves to verbalize the table. However, the quality of the generated text is not optimal due to the challenges of open-ended text generation. Several existing studies, such as Table-to-Text [51], have explored this sub-task. In the future, we can enhance performance and efficiency by replacing this step with specifically trained small language models, which could further improve the semantic density of the verbalized table.
@@ Appendix D Detailed Settings of Challenge Analysis Experiments
We conduct extensive experiments to analyze the challenges of table understanding with language models (LMs). Specifically, we perform challenge analysis experiments on the WikiTQ dataset [11], which consists of 4,344 data instances. Following previous work [17, 46], we use the exact match of the final answer as the evaluation metric to measure accuracy. Our experiments utilize OpenAI models hosted on Microsoft Azure. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. For each model, we use the following versions: gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), and o1 (o1-preview-0912).
@@
Appendix E Extended Experiments on Additional Table-Understanding Benchmarks
We perform additional experiments on diverse table-understanding tasks to further assess the robustness of TableMaster.
@@ E.1 Evaluation on Free-form QA with the FetaQA Dataset
Fine-Tuning (T5-large) [16] 30.54 63 41 53
End-to-End QA (Codex) [64] 27.96 62 40 52
End-to-End QA (PaLM 2) [17] 28.37 63 41 53
Dater (PaLM 2) [16] 29.47 63 41 53
Chain-of-Table (PaLM 2) [17] 32.61 (+14.9%) 66 (+4.8%) 44 (+7.3%) 56 (+5.7%)
@@
PaLM 2 has been deprecated [65] and is no longer accessible. Therefore, we use a comparable language model, gpt-4o, to conduct experiments on FetaQA and compare the results with previous methods. Additionally, we use 20 exemplars for few-shot in-context learning to align with the dataset’s format.
@@
However, the improvement of TableMaster over baseline methods remains marginal, with some values even falling below those of previous approaches in absolute terms. We believe this does not fully reflect the model’s actual performance in free-form QA. We attribute this to the n-gram text similarity metrics used in ROUGE-1/2/L [59], which are known to be insensitive to improvements gained from in-context learning [66]. These metrics struggle to capture stylistic and structural enhancements in free-form text generation. Since models rely on instructions and a limited number of examples, they may not fully adapt to the expected output format, leading to an underestimation of performance gains.
@@ E.2 Evaluation on Hierarchical Tables with the HiTab Dataset
UUnlike WikiTQ and TabFact, the HiTab dataset [60] contains hierarchical tables that violate the flat row–column assumption. Such structure challenges models to reason across multi-level headers and parent–child cell relations.
We employ a GPT-4o backbone and follow HiTab’s official data split. MultiCoT [67] extends Chain-of-Table reasoning to multiple tables; both MultiCoT and TableMaster operate on the same converted relational tables. E5 [68] represents the current SOTA on HiTab, being explicitly tailored for complex tables. Because our framework (TableMaster) is designed for relational tables, we introduce a lightweight relational-table converter via prompting o1: each hierarchical table is decomposed into several relational subtables, while contextual tags are propagated to preserve structural cues.
@@
   E5 [68] 77.3
Table 4 shows that TableMaster outperforms the strongest chain-of-table baseline by +0.7 point (from 73.5 to 74.2). Residual errors are mainly due to information loss during the conversion step. Future work will integrate hierarchical relations directly into the reasoning module.
@@ E.3 Evaluation on Numerical Reasoning with the FinQA Dataset
FinQA [61] requires multi-step numerical reasoning over financial reports—e.g., computing growth rates or combining multiple cells with arithmetic operators. Hence it evaluates whether TableMaster can execute numerical formulas correctly, not just extract text spans. We keep the same training recipe as in. Two backbones are considered: GPT-4o-mini (4m) and GPT-4o (4o).
Table 5: Accuracy on FinQA. TableMaster consistently boosts numerical-reasoning accuracy over both backbones.
@@ Appendix F Table Understanding Baselines
To better facilitate future research, we evaluate different reasoning methods across various base models. Table 6 presents the accuracy results of our reproduced baselines on WikiTQ and TabFact, comparing different base LLMs and reasoning methods. The table includes evaluations on o1-preview ($\sim$300B), o1-mini ($\sim$100B), gpt-4o ($\sim$200B), gpt-4o-mini ($\sim$8B), and gpt-3.5-turbo ($\sim$175B). Each model is tested with various reasoning strategies, including Direct, chain of thought, and Program of Thought, alongside our proposed TableMaster.
@@
The performance gap is even more pronounced for gpt-3.5-turbo, where TableMaster reaches 68.21% on WikiTQ and 83.65% on TabFact, significantly outperforming both chain of thought (59.92%, 69.52%) and Program of Thought (50.32%, 68.82%). Interestingly, we observe that while TableMaster ’s improvement is limited on gpt-4o, the weaker the base model, the greater the performance improvement. While o1-preview and o1-mini achieve high accuracy with the Direct method (84.60%, 92.05% for o1-preview and 83.49%, 91.35% for o1-mini), the results of TableMaster on gpt-4o demonstrate that our method is capable of achieving state-of-the-art performance across different LL architectures.
Additionally, we find that chain of thought reasoning is highly effective, achieving strong accuracy across models. Even a simple chain of thought approach outperforms previous methods that rely solely on symbolic reasoning [41], indicating that chain of thought should be retained as a key component in the reasoning framework.
@@ Appendix G Performance Analysis Under Different Table Sizes
Binder [18] 56.54 26.13 6.41
Dater [16] 62.50 42.34 34.62
Chain-of-Table [17] 68.13 52.25 44.87
@@
Table 7 presents a performance comparison across different table sizes, categorized into small (<2k tokens), medium (2$\sim$4k tokens), and large (>4k tokens). The results compare several methods, including Binder [18], Dater [16], and Chain-of-Table [17], against TableMaster. All methods are evaluated using gpt-3.5-turbo, with additional results of TableMaster provided for gpt-4o-mini.
@@ I.1 Theoretical Analysis
Total $\text{Total Cost}=(2\,k+1)\times n+(e+2.5)\times(a\times b).$ (1)
@@ I.2 Empirical Analysis
Recent advancements in table understanding, such as Chain-of-Table [17] and Tree-of-Table [45], involve a step-by-step evolution of tables through a long chain of transformations. In each new step, both the original table and the newly generated sub-table must be processed by language models. Additionally, their iterative process is complex, unstable, and difficult to analyze theoretically. In contrast, our approach is general and comprehensive, avoiding the trivial overhead of sub-table extraction. Instead, it focuses on holistic reasoning while maintaining ideal efficiency.
@@
0:  $C^{0}$: Initially selected columns
0:  $\mathbb{C}$: Ranked candidate column indices
0:  $\mathbb{T}^{F}$: Final table-of-focus
@@
Self-consistency methods, which perform multiple inferences to improve reliability, achieve better results. Five-shot CoT self-consistency reaches 74.98%, while five-shot PoT self-consistency lags behind at 63.97%. As introduced in [46], mixed self-consistency (3 CoT + 3 PoT) and (5+5) further improve accuracy to 76.70% and 77.46%, respectively, at the cost of increased inference time (6 and 10 passes). Self-evaluation (self-eval) first performs CoT and PoT inferences (Prompt 26), then selects the better result, achieving 70.58% with 2 inferences.
@@
PoTable (Previous SOTA) [41] 64.73 88.93
@@
To highlight the performance of TableMaster, we report the best scores of 78.13 and 90.12 in the main results table. Regardless, our method consistently outperforms the previous state-of-the-art, PoTable [41], on both WikiTQ and TabFact.

### v4

@@ 5.3 Ablation Study
Reasoning. The reasoning stage exhibits the most significant performance drop when removed. Removing textual reasoning leads to the largest accuracy decline, with a 4.28% drop on WikiTQ and 1.73% on TabFact, underscoring its necessity for complex reasoning tasks. Similarly, removing symbolic reasoning results in a 2.03% and 0.79% drop on WikiTQ and TabFact, respectively, demonstrating that symbolic reasoning enhances numerical and structured table interpretations. Finally, removing textual guidance, which improves the semantic flexibility of symbolic reasoning, reduces accuracy by 2.92% on WikiTQ and 0.44% on TabFact. This highlights that textual guidance is particularly beneficial and important in symbolic reasoning by ensuring alignment with the problem context. More analysis of adaptive reasoning can be found at Appendix L.
@@ 6 Conclusion
In this paper, we explore table understanding with language models. Given the characteristics of tabular data, we identify key challenges in table understanding. To overcome these challenges, we propose TableMaster, a recipe and comprehensive framework that integrates multiple solutions. Extensive analyses and experiments demonstrate our findings and the effectiveness of TableMaster. In the future, we plan to extend and refine the framework to improve its performance across diverse practical applications, where discussed in Appendix B.
Appendix A Ethics Statement
TableMaster introduces a general-purpose, modular framework that improves the ability of language models (LMs) to understand and reason over tabular data. Its applications span domains such as business intelligence, scientific reporting, education, and healthcare, where structured data plays a critical role. By enabling adaptive reasoning across textual and symbolic paradigms, TableMaster improves both accuracy and transparency in question answering, verification, and analysis over tables. These advances may lead to better decision-support systems and streamlined human–AI collaboration in spreadsheet-heavy workflows.
Despite these benefits, there are potential risks. As TableMaster automates reasoning over tabular content, it could be misused to generate misleading analyses or automate decisions without adequate human oversight. Furthermore, inaccurate reasoning, especially when symbolic operations are applied incorrectly, could result in flawed conclusions or financial misjudgments. Biases in training data might also manifest in generated answers or program logic. In addition, as the framework relies on LLMs’ capabilities, disparities across languages, domains, or spreadsheet conventions may lead to uneven performance, potentially disadvantaging users in low-resource settings.
To mitigate these concerns, we propose several safeguards. First, outputs involving symbolic reasoning should be verified via deterministic execution (e.g., code validation or unit tests) before downstream use. Second, we encourage model evaluation on a diverse range of real-world tables, including messy, hierarchical, or multilingual formats. Third, human oversight is recommended in high-stakes applications, especially when deployed in financial, legal, or healthcare settings. Fourth, interpretability tools—such as reasoning traces or program annotations—should be integrated to facilitate debugging and auditing. Finally, we advocate for transparent reporting of model limitations and publishing benchmark results across different domains and table types to promote responsible deployment.
Appendix B Limitations, Extendability, and Future Works
@@
B.1 Technical Refinement
Wild Table. In our experiments, the tables in the three datasets we use are already cleaned; therefore, we do not explicitly implement table normalization in our evaluation experiments. However, we conduct analysis experiments to highlight the importance of table normalization for handling wild tables. In practical scenarios, various tools are available for table normalization. Regular expression matching can be employed for formatting, and small language models can also be leveraged to efficiently process and normalize tables (Nahid and Rafiei, 2024a).
@@ B.1 Technical Refinement
Table-of-Focus Construction. In designing the Table-of-Focus, we employ two efficient methods: LM prompting for column lookup and SQL generation for row lookup. The Table-of-Focus is then constructed based on the results of these two lookups. Many previous works (Ji et al., 2024; Wang et al., 2024) have introduced complex approaches for extracting relevant sub-tables. In contrast, our method remains intentionally simple, prioritizing efficiency and adaptability. We believe that in the future, more advanced techniques may emerge to further enhance the extraction of key information.
Table Verbalization. To facilitate the implementation of TableMaster, we utilize language models themselves to verbalize the table. However, the quality of the generated text is not optimal due to the challenges of open-ended text generation. Several existing studies, such as Table-to-Text (Parikh et al., 2020), have explored this sub-task. In the future, we can enhance performance and efficiency by replacing this step with specifically trained small language models, which could further improve the semantic density of the verbalized table.
@@ B.2 Downstream Applications
Appendix C Datasets Used for Evaluation
Table 3 shows all datasets use for evluation in this study, license and source are also included.
Table 3: Benchmarks used for evaluation.
Dataset # Test Table Type Domain License Source
WikiTQ (Pasupat and Liang, 2015) 4,217 Relational Wikipedia $C$-BY-SA-4.0 Link
TabFact (Chen et al., 2020) 2,024 Relational Wikipedia $C$-BY-4.0 Link
FetaQA (Nan et al., 2022) 1,165 Relational Wikipedia $C$-BY-SA-4.0 Link
FinQA (Chen et al., 2021b) 1,147 Relational Finance MIT Link
HiTab (Cheng et al., 2022) 1,583 Hierarchical Reports C-UDA 1.0 Link
@@ Appendix D Detailed Settings of Challenge Analysis Experiments
We conduct extensive experiments to analyze the challenges of table understanding with language models (LMs). Specifically, we perform challenge analysis experiments on the WikiTQ dataset (Pasupat and Liang, 2015), which consists of 4,344 data instances. Following previous work (Wang et al., 2024; Liu et al., 2024b), we use the exact match of the final answer as the evaluation metric to measure accuracy. Our experiments utilize OpenAI models hosted on Microsoft Azure [^1] https://azure.microsoft.com/en-us/support/legal/. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. For each model, we use the following versions: gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), and o1 (o1-preview-0912).
@@
Appendix E Extended Experiments on Additional Table Understanding Benchmarks
We perform additional experiments on diverse table‑understanding tasks to further assess the robustness of TableMaster.
@@ E.1 Evaluation on Free-form QA with the FetaQA Dataset
Fine-Tuning (T5-large) (Ye et al., 2023) 30.54 63 41 53
End-to-End QA (Codex) (Chen et al., 2021a) 27.96 62 40 52
End-to-End QA (PaLM 2) (Wang et al., 2024) 28.37 63 41 53
Dater (PaLM 2) (Ye et al., 2023) 29.47 63 41 53
Chain-of-Table (PaLM 2) (Wang et al., 2024) 32.61 (+14.9%) 66 (+4.8%) 44 (+7.3%) 56 (+5.7%)
@@
PaLM 2 has been deprecated (Anil et al., 2023) and is no longer accessible. Therefore, we use a comparable language model, gpt-4o, to conduct experiments on FetaQA and compare the results with previous methods. Additionally, we use 20 exemplars for few-shot in-context learning to align with the dataset’s format.
@@
However, the improvement of TableMaster over baseline methods remains marginal, with some values even falling below those of previous approaches in absolute terms. We believe this does not fully reflect the model’s actual performance in free-form QA. We attribute this to the n-gram text similarity metrics used in ROUGE-1/2/L (Lin, 2004), which are known to be insensitive to improvements gained from in-context learning (Maynez et al., 2023). These metrics struggle to capture stylistic and structural enhancements in free-form text generation. Since models rely on instructions and a limited number of examples, they may not fully adapt to the expected output format, leading to an underestimation of performance gains.
@@ E.2 Evaluation on Hierarchical Tables with the HiTab Dataset
UUnlike WikiTQ and TabFact, the HiTab dataset (Cheng et al., 2022) contains hierarchical tables that violate the flat row–column assumption. Such structure challenges models to reason across multi‑level headers and parent–child cell relations.
We employ a GPT‑4o backbone and follow HiTab’s official data split. MultiCoT (CYQIQ, 2025) extends Chain‑of‑Table reasoning to multiple tables; both MultiCoT and TableMaster operate on the same converted relational tables. E5 (Zhang et al., 2024c) represents the current SOTA on HiTab, being explicitly tailored for complex tables. Because our framework (TableMaster) is designed for relational tables, we introduce a lightweight relational‑table converter via prompting o1: each hierarchical table is decomposed into several relational subtables, while contextual tags are propagated to preserve structural cues.
@@
   E5 (Zhang et al., 2024c) 77.3
Table 5 shows that TableMaster outperforms the strongest chain‑of‑table baseline by +0.7 point (from 73.5 to 74.2). Residual errors are mainly due to information loss during the conversion step. Future work will integrate hierarchical relations directly into the reasoning module.
@@ E.3 Evaluation on Numerical Reasoning with the FinQA Dataset
FinQA (Chen et al., 2021b) requires multi‑step numerical reasoning over financial reports—e.g., computing growth rates or combining multiple cells with arithmetic operators. Hence it evaluates whether TableMaster can execute numerical formulas correctly, not just extract text spans. We keep the same training recipe as in. Two backbones are considered: GPT‑4o‑mini (4m) and GPT‑4o (4o).
Table 6: Accuracy on FinQA. TableMaster consistently boosts numerical‑reasoning accuracy over both backbones.
@@ Appendix F Table Understanding Baselines
To better facilitate future research, we evaluate different reasoning methods across various base models. Table 7 presents the accuracy results of our reproduced baselines on WikiTQ and TabFact, comparing different base LLMs and reasoning methods. The table includes evaluations on o1-preview ($\sim$300B), o1-mini ($\sim$100B), gpt-4o ($\sim$200B), gpt-4o-mini ($\sim$8B), and gpt-3.5-turbo ($\sim$175B). The exact number of parameters for several LMs (e.g., GPT, o1) has not been publicly disclosed. Most parameter counts are estimates reported to provide context for understanding model performance. For more precise information, please refer to the original or future official documentation (Abacha et al., 2025). Each model is tested with various reasoning strategies, including Direct, chain of thought, and Program of Thought, alongside our proposed TableMaster.
@@
The performance gap is even more pronounced for gpt-3.5-turbo, where TableMaster reaches 68.21% on WikiTQ and 83.65% on TabFact, significantly outperforming both chain of thought (59.92%, 69.52%) and Program of Thought (50.32%, 68.82%). Interestingly, we observe that while TableMaster ’s improvement is limited on gpt-4o, the weaker the base model, the greater the performance improvement. While o1-preview and o1-mini achieve high accuracy with the Direct method (84.60%, 92.05% for o1-preview and 83.49%, 91.35% for o1-mini), the results of TableMaster on gpt-4o demonstrate that our method is capable of achieving state-of-the-art performance across different $L$ architectures.
Additionally, we find that chain of thought reasoning is highly effective, achieving strong accuracy across models. Even a simple chain of thought approach outperforms previous methods that rely solely on symbolic reasoning (Mao et al., 2024), indicating that chain of thought should be retained as a key component in the reasoning framework.
@@ Appendix G Performance Analysis Under Different Table Sizes
Binder (Cheng et al., 2023) 56.54 26.13 6.41
Dater (Ye et al., 2023) 62.50 42.34 34.62
Chain-of-Table (Wang et al., 2024) 68.13 52.25 44.87
@@
[…]
@@ I.1 Theoretical Analysis
$\text{Total Cost}=(2\,k+1)\times n+(e+2.5)\times(a\times b).$ (1)
@@ I.2 Empirical Analysis
Recent advancements in table understanding, such as Chain-of-Table (Wang et al., 2024) and Tree-of-Table (Ji et al., 2024), involve a step-by-step evolution of tables through a long chain of transformations. In each new step, both the original table and the newly generated sub-table must be processed by language models. Additionally, their iterative process is complex, unstable, and difficult to analyze theoretically. In contrast, our approach is general and comprehensive, avoiding the trivial overhead of sub-table extraction. Instead, it focuses on holistic reasoning while maintaining ideal efficiency.
On the first 100 examples of TabFact, we evaluate TableMaster against the representative baseline Chain-of-Table (Wang et al., 2024) using GPT-4o, ensuring a fair comparison without self-consistency decoding.
Table 9: Token usage comparison on the TabFact subset using GPT-4o.
Method Prompt Tokens Completion Tokens
Chain-of-Table 13,209.6 914.2
TableMaster (Ours) 3,393.5 738.6
As shown in Table 9, TableMaster consumes substantially fewer tokens while maintaining strong performance. While our design may appear intricate, it integrates practical features such as fallbacks to early-exit or full-table reasoning, thereby avoiding unnecessary table transformations. In contrast, Chain-of-Table continuously evolves the table without fallback safeguards, leading to much higher token usage. This limitation is not acknowledged in the original paper, whereas our framework explicitly incorporates token efficiency as a design principle.
@@
[…]
@@
We consider adaptive reasoning a key component in table understanding. Concurrent work, such as Abhyankar et al. (2025), also explores this direction.
@@
Self-consistency methods, which perform multiple inferences to improve reliability, achieve better results. Five-shot CoT self-consistency reaches 74.98%, while five-shot PoT self-consistency lags behind at 63.97%. As introduced in (Liu et al., 2024b), mixed self-consistency (3 CoT + 3 PoT) and (5+5) further improve accuracy to 76.70% and 77.46%, respectively, at the cost of increased inference time (6 and 10 passes). Self-evaluation (self-eval) first performs CoT and PoT inferences (Prompt 26), then selects the better result, achieving 70.58% with 2 inferences.
@@
PoTable (Previous SOTA) (Mao et al., 2024) 64.73 88.93
@@
To highlight the performance of TableMaster, we report the best scores of 78.13 and 90.12 in the main results table. Regardless, our method consistently outperforms the previous state-of-the-art, PoTable (Mao et al., 2024), on both WikiTQ and TabFact.
@@
Appendix R The Use of Large Language Models (LLMs)
In this work, large language models (LLMs) were used only to aid with writing and polishing the manuscript. Specifically, LLMs were employed for grammar correction, phrasing suggestions, and improving readability. All research ideas, methodological contributions, theoretical analyses, and experiments were entirely conceived, designed, and executed by the authors without the involvement of LLMs. The authors take full responsibility for the scientific content of the paper.